OpenStudy (anonymous):

I need some help with writing an equation of a line thats parallel to y=2x-7 and passes through the point (3,4)

OpenStudy (spareb665):

y=2x-2

OpenStudy (rosedewittbukater):

To get a parallel graph, you need to have the same slope m as the original equation. So the new equation would be y=2x+b. Plug in the point to find a line that passes through it. \[4=2(3)+b\] Next you have to find b. \[4=6+b\] \[-2=b\] The final equation will be \[y=2x-2\]

OpenStudy (mathstudent55):

A line that has a given slope, m, and passes through point (x1, y1) has the equation: y = y1 = m(x - x1) You are given a point, (3, 4) which is (x1, y1), so you have x1 = 3, and y1 = 4. All you need is the slope of the parallel line. Parallel lines have the same slope. So the slope you need is the same as the slope of y = 2x - 7 y = 2x - 7 is written in the slope-intercept form, which is y = mx + b, where m = slope, and b = y-intercept Just by looking at y = 2x - 7, you can see that the slope is m = 7. Now use m = 7, x1 = 3, and y1 = 4 in the equation of a line above and find the equation of the line.

OpenStudy (anonymous):

thanks this really helped me out guys